OpenStudy (anonymous):

What is the exact circumference of the circle? A. 16 cm B. 8 cm C. 4 cm D. 2 cm

OpenStudy (anonymous):

The formula for the circumference of a circle is 2 x \(\Huge{\pi}\) x radius

OpenStudy (michele_laino):

Please , note that you have to apply this formula: C=2*pi*r, where pi=3.14.., and r is the radius of your circumference

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, @zahviem does it tell you the radius?